OpenStudy (theviper):

Please help:) The H.C.F. of two expressions is x & their L.C.M is\[\large{x^3-9x}\]If one of the expressions is \[\large{x^2+3x}\], then the other expression is??? @jim_thompson5910 @experimentX :)

OpenStudy (experimentx):

\[ x^3 - 9x = x(x^2 - 3^2) = x(x+3)(x - 3) \\ x^2 + 3x = x(x+3)\]

OpenStudy (theviper):

It's \[x^2-3x\]in my book:\

OpenStudy (theviper):

@experimentX

OpenStudy (theviper):

I don't want to disturb but @satellite73

jimthompson5910 (jim_thompson5910):

Hint: LCM = [ (Expression 1)*(Expression 2) ]/(HCF)

OpenStudy (anonymous):

what is missing from \(x(x+3)\) to get \(x(x+3)(x-3)\)?

OpenStudy (theviper):

(x-3)

OpenStudy (anonymous):

yup

OpenStudy (theviper):

k thanx everyone i got it:)

OpenStudy (theviper):

but i got @experimentX u r right thanx:)

OpenStudy (experimentx):

yw